# Spatial Impulse Response Analysis and Ensemble Learning for Efficient Precision Level Sensing

### v1

@@ Abstract
In this paper, we propose an innovative method for determining the fill level of containers, such as trash cans, addressing a critical aspect of waste management. The method combines spatial impulse response analysis with machine learning techniques, offering a unique and effective approach for sound-based classification that can be extended to various domains beyond waste management. By employing a buzzer-generated sine sweep signal, we create a distinctive signature specific to the fill level of the waste container. This signature is then interpreted by a specially developed ensemble learning algorithm. Our approach achieves a classification accuracy of over 90% when implemented locally on a development board, eliminating the need to delegate complex classification tasks to external entities. Using low-cost and energy-efficient hardware components, our method offers a cost-effective approach that contributes to sustainable and efficient waste management practices, providing a reliable and locally deployable solution.
{IEEEkeywords} Artificial intelligence (AI), machine learning (ML), spatial impulse response, waste management.
@@ 1 Introduction
    Achieve local sound classification with reliable accuracy, on an ESP8266 microcontroller. This aspect of our work is particularly noteworthy as it merges high-end performance with practical applicability, ensuring our methodology is both efficient and accessible.
@@ 1.2 System Overview
This project focuses on the development of an IWMS that utilizes on-device AI for the classification of waste container fill levels. As a proof of concept, the inner side of a waste container’s lid was equipped with an ESP8266, a compact Wi-Fi-capable micro-controller [19], to which a passive buzzer and a microphone were attached (Figure 1). The ESP8266 served as the central processing unit of the system, generating sound signals through the passive buzzer and capturing the resulting acoustic response via the microphone. The captured signal was then transmitted to a Raspberry Pi for storage using MQTT [20], a lightweight messaging protocol designed for small sensors and mobile devices.
The primary objective of this research is to accurately classify the fill levels of waste containers into five distinct categories: empty (0%), quarter-filled (25%), half-filled (50%), three-quarters filled (75%), and completely filled (100%). To simulate waste materials, straw and cardboard were chosen due to their varying acoustic response characteristics. Central to our research is the SIREC algorithm, which is based on ensemble learning principles. Ensemble learning algorithms are well-established in the field of machine learning and are known for their ability to generate robust and accurate predictions by combining the decisions of multiple models. In this study, we employ an ensemble of decision tree models to classify the fill levels. Decision trees were selected for their versatility in handling complex datasets and their ability to provide transparent and interpretable decision pathways.
@@ 2.2 Training & Testing
Following the extraction of RIRs, the data is transmitted to a Raspberry Pi via MQTT. There, Python scripts manage and utilize the data for two primary processes, depicted as numbers 2 and 3 in Figure 2, which involve the SIREC algorithm developed for this project. During the training process, RIRs corresponding to varying fill levels of the waste container are collected and labeled. An ensemble ML model, specifically the SIREC model, is then trained using these RIRs, employing a predetermined number of estimators ($N_{\text{estim.}}$), which are decision trees. For each tree, two intervals are generated: a random interval $I_{\text{rnd.}}$, whose length varies within a adjustable maximum and minimum range across the total length of the RIRs, and a fixed interval $I_{\text{fix.}}$, matching the length of the random interval but positioned at the start of the RIRs. Features are then extracted from both intervals to train the respective tree. Within $I_{\text{rnd.}}$, the magnitude of the frequency spectrum is determined using FFT, and the ratio of the minimum to maximum values is extracted. Empirical analysis has shown these features to be significant for classification. Within $I_{\text{fix.}}$, two feature classes are also extracted, but based on novel approaches, incorporating the mean and standard deviation as characteristic features of the time series into the tree’s training. Time series data can have varying structures, and in some cases, they may exhibit symmetry along the x-axis without a vertical shift. In this project, the RIR is calculated in a manner that closely exhibits these two properties. An example of an extracted RIR is shown in Figure 4. Considering this structure of the RIRs, calculating the mean would yield a value close to zero for different classes, making it also problematic for the standard deviation, which depends on the mean. In light of this, we propose an alternative calculation method for the fixed interval, computing the difference between each pair of adjacent values across all $N-1$ values to determine the mean and standard deviation of these newly generated $N-1$ values. This approach captures not only the magnitude of the mean and standard deviation but also the characteristic property of each RIR, combined with these two measures. The necessary computational steps are outlined in Equations 2 and 3.
@@
Each of the $N_{\text{estim.}}$ trees is trained using this feature extraction methodology, forming an independent ML model that, in combination, constitutes SIREC, the complete ensemble ML model. In the testing process, depicted as number 3 in Figure 2, the model’s accuracy is validated using a portion of the collected RIRs and their corresponding fill level labels. Here, SIREC utilizes the intervals created during its training process, extracts relevant features, and aggregates predictions from all its decision trees to derive a final result. The F1 score metric is then used to determine if the validation is sufficiently accurate to, if successful, convert the trained and validated SIREC model into C++ code. This code can then be seamlessly uploaded to the ESP8266 board using the Arduino IDE [31]–a development environment that facilitates programming and interfacing with various micro-controllers and modules. Thus, enabling local predictions using the trained ML model without the need to delegate the classification task to external resources.
@@ 3 Validation
Resources from our research are accessible at https://gitlab.rlp.net/rgdsai/sirec/.
@@ 3.1 Performance
Figure 5 presents the classification accuracy of various algorithms, evaluated based on the F1 score metric using a segment length of 300, where each score represents the mean accuracy derived from 100 iterations. It is important to note that, throughout this paper, the term ”segment length” refers to the length of a subsection extracted from the total duration of the RIRs. The algorithms presented in Figure 5, with the exception of SIREC which has been previously discussed, are selected from the sktime [32] and sklearn [33] libraries, which are widely used in ML for time series and general-purpose classification tasks, respectively. Notably, all ensemble-based algorithms in this comparison were configured with an equal number of estimators (100) to ensure a consistent basis for comparison. Furthermore, Figure 6 illustrates one performance of SIREC via a confusion matrix, which is a tabular representation of the model’s correct and incorrect predictions across different classes. This matrix’s rows correspond to actual classes, while the columns indicate the predicted classes by the model. The entries on the main diagonal (from top-left to bottom-right) represent the number of correct predictions, aligning with the true positives for each class. All other entries denote misclassifications.
Figure 5: Mean F1 scores of different classification models (100 iterations). Training data: $T_{1}$ $T_{2}$, Test data: $T_{3}$ $T_{4}$. Inspected sample segment length: 300.
Figure 6: Example of a confusion matrix for SIREC.
It is particularly noteworthy that SIREC and the Random Interval Spectral Ensemble (RISE) [34] algorithm, which secured second place in Figure 5, share two fundamental commonalities. Both methodologies are predicated on an ensemble of decision trees and both engage in the generation of random intervals, within which they extract the magnitude of the frequency spectrum as one of the features. However, there are key advantages that distinguish SIREC from RISE, rendering it more beneficial. Contrary to RISE, SIREC exhibits superior speed and energy efficiency (as will be discussed in the following section and shown in Table 3), as well as a significant improvement in memory efficiency. This is largely due to the fact that RISE computes the auto-correlation function, with the default setting being 100 auto-correlation terms, which inherently makes SIREC more suited for local operation on micro-controllers for classification tasks. Furthermore, SIREC tends to demonstrate marginally better average accuracy. To support this claim, Figure 7, as an extension to Figure 5, provides a detailed depiction of the F1 scores between RISE and SIREC over 100 iterations, where both algorithms employ an estimator count of 100, with all other parameters retained at their default settings.
@@
A significant advantage of the Random Interval Generation concept for feature extraction, as implemented in SIREC, is its inherent reduction or elimination of the need for expert knowledge about the collected data to train the model. For instance, a stochastic search across various randomly generated intervals within specified minimum and maximum lengths allows for the bypassing of such expert knowledge. This stochastic search method ensures the identification of the ideal data segment for feature extraction without necessitating extensive manual data analysis. Moreover, since SIREC is based on decision trees, which represent a symbolic form of representation, acquiring expert knowledge about the data post-training is straightforward. Our stochastic search for optimal SIREC parameters (thus, the best random intervals) has achieved accuracies up to 100% in some cases. The results of this stochastic search, along with the parameters used for SIREC, are presented in Table 2.
@@ 3.2 Energy measurements
In the analysis of energy consumption, our investigation reveals significant insights into the trade-offs between computational energy efficiency and predictive accuracy across three distinct algorithms: our proposed algorithm (SIREC), Time Series Forest Classifier (TSFC)[35] and RISE. The measurement method and experiment setup is based upon our previous work [36, 37] and explained in detail by [38]. The resource measurements adhere to the methodology and guidelines outlined in the Green Software Measurement Model [39]. Measurements were carried out using a computer system that included an Intel Core i5-650 CPU. The system’s memory was composed of two 2 GB RAM modules, totaling 4 GB. For storage, the computer utilized a dual-setup, comprising a 500 GB hard disk drive (HDD) and a 250 GB solid-state drive (SSD). The analysis is structured around various scenarios–configurations of segment lengths and numbers of estimators, focusing on key metrics such as energy consumption during training and testing phases, computational time, CPU usage, and F1 score. These scenarios, along with their corresponding results listed in Table 3, were tested on the same datasets, T1 and T2 for training, and T3 and T4 for testing, as detailed in Table 1. It is important to highlight that a random state parameter was employed across all algorithms to ensure comparability and control over experimental variability. A value of 7 was selected arbitrarily for the random state, with no intent to favor any specific algorithm, especially not SIREC.

### v2

@@ Abstract
In this paper, we propose an innovative method for determining the fill level of containers, such as trash cans, addressing a critical aspect of waste management. The method combines spatial impulse response analysis with machine learning (ML) techniques, offering a unique and effective approach for sound-based classification that can be extended to various domains beyond waste management. By employing a buzzer-generated sine sweep signal, we create a distinctive signature specific to the fill level of the waste container. This signature, once accurately decoded, is then interpreted by a specially developed ensemble learning algorithm. Our approach achieves a classification accuracy of over 90% when implemented locally on a development board, optimizing operational efficiencies and eliminating the need to delegate complex classification tasks to external entities. Using low-cost and energy-efficient hardware components, our method offers a cost-effective approach that contributes to sustainable and efficient waste management practices, providing a reliable and locally deployable solution.
{IEEEkeywords}
Artificial intelligence (AI), level sensing, machine learning (ML), smart waste bin, sound classification, spatial impulse response, waste management.
@@ 1 INTRODUCTION
    Achieve local sound classification with reliable accuracy, on an ESP8266 microcontroller. This aspect of our work is particularly noteworthy as it merges solid performance with practical applicability, ensuring our methodology is both efficient and accessible.
@@ 1.2 SYSTEM OVERVIEW
This project focuses on the development of an IWMS that utilizes on-device AI for the classification of waste container fill levels. As a proof of concept, the inner side of a waste container’s lid was equipped with an ESP8266, a compact Wi-Fi-capable micro-controller [19], to which a passive buzzer and a microphone were attached, as shown in Figure 1. The ESP8266 board (4) served as the central processing unit of the system, generating sound signals (sine sweeps) through the passive buzzer (2) and capturing the resulting acoustic response (bin interior) via the microphone (3). The captured and preprocessed (see Section 2) signal was then transmitted to a Raspberry Pi for storage using MQTT [20], a lightweight messaging protocol designed for small sensors and mobile devices.
The primary objective of this research is to accurately classify the fill levels of waste containers into five distinct categories: empty (0%), quarter-filled (25%), half-filled (50%), three-quarters filled (75%), and completely filled (100%). To simulate waste materials, straw and cardboard were chosen due to their varying acoustic response characteristics. Central to our research is the SIREC algorithm, which is based on ensemble learning principles. Ensemble learning algorithms are well-established in the field of ML and are known for their ability to generate robust and accurate predictions by combining the decisions of multiple models. In this study, we employ an ensemble of decision tree (DT) models to classify the fill levels. DTs were selected for their versatility in handling complex datasets and their ability to provide transparent and interpretable decision pathways.
@@ 2.2 TRAINING & TESTING
Following the extraction of RIRs, the data is transmitted to a Raspberry Pi via MQTT. There, Python scripts manage and utilize the data for two primary processes, depicted as numbers 2 and 3 in Figure 2, which involve the SIREC algorithm developed for this project. During the training process, RIRs corresponding to varying fill levels of the waste container are collected and labeled. An ensemble ML model, specifically the SIREC model, is then trained using these RIRs, employing a predetermined number of estimators ($N_{\text{estim.}}$), which are DTs. For each tree, two intervals are generated: a random interval $I_{\text{rnd.}}$, whose length varies within a adjustable maximum and minimum range across the total length of the RIRs, and a fixed interval $I_{\text{fix.}}$, matching the length of the random interval but positioned at the start of the RIRs. Features are then extracted from both intervals to train the respective tree. Within $I_{\text{rnd.}}$, the magnitude of the frequency spectrum is determined using FFT, and the ratio of the minimum to maximum values is extracted. Empirical analysis has shown these features to be significant for classification. Within $I_{\text{fix.}}$, two feature classes are also extracted, but based on novel approaches, incorporating the mean and standard deviation as characteristic features of the time series into the tree’s training. Time series data can have varying structures, and in some cases, they may exhibit symmetry along the x-axis without a vertical shift. In this project, the RIR is calculated in a manner that closely exhibits these two properties. An example of an extracted RIR is shown in Figure 4. Considering this structure of the RIRs, calculating the mean would yield a value close to zero for different classes, making it also problematic for the standard deviation, which depends on the mean. In light of this, we propose an alternative calculation method for the fixed interval, computing the difference between each pair of adjacent values across all $N-1$ values to determine the mean and standard deviation of these newly generated $N-1$ values. This approach captures not only the magnitude of the mean and standard deviation but also the characteristic property of each RIR, combined with these two measures. The necessary computational steps are outlined in Equations 2 and 3.
@@
Each of the $N_{\text{estim.}}$ trees is trained using this feature extraction methodology, forming an independent ML model that, in combination, constitutes SIREC, the complete ensemble ML model. In the testing process, depicted as number 3 in Figure 2, the model’s accuracy is validated using a portion of the collected RIRs and their corresponding fill level labels. Here, SIREC utilizes the intervals created during its training process, extracts relevant features, and aggregates predictions from all its DTs to derive a final result. The F1 score metric is then used to determine if the validation is sufficiently accurate to, if successful, convert the trained and validated SIREC model into C++ code. This code can then be seamlessly uploaded to the ESP8266 board using the Arduino IDE [31]–a development environment that facilitates programming and interfacing with various micro-controllers and modules. Thus, enabling local predictions using the trained ML model without the need to delegate the classification task to external resources.
@@ 3.1 PERFORMANCE
Figure 5 displays the classification accuracy results of various ML models, represented through box plots for the F1 score metric and confusion matrices normalized with respect to the true labels. The results are based on 100 iterations of training and testing, considering a segment length of 300. It is important to note that, throughout this paper, the term ”segment length” refers to the length of a subsection extracted from the total duration of the RIRs. The algorithms presented in Figure 5, with the exception of SIREC which has been previously discussed, are selected from the sktime [32] and sklearn [33] libraries, which are widely used in ML for time series and general-purpose classification tasks, respectively. Notably, all ensemble-based algorithms in this comparison were configured with an equal number of estimators (100), keeping all other parameters at their default values to ensure a consistent basis for comparison.
Figure 5: Box plots (F1 score; red dashed line indicating the mean; solid line indicating the median) and confusion matrices (normalized with respect to true labels) for the considered ML models over 100 iterations of training and testing. Training data: $T_{1}$ $T_{2}$, Test data: $T_{3}$ $T_{4}$. Sample segment length: 300.
It is particularly noteworthy that SIREC and the Random Interval Spectral Ensemble (RISE) [34] algorithm, which secured second place in Figure 5, share two fundamental commonalities. Both methodologies are predicated on an ensemble of DTs and both engage in the generation of random intervals, within which they extract the magnitude of the frequency spectrum as one of the features. However, there are key advantages that distinguish SIREC from RISE, rendering it more beneficial. Contrary to RISE, SIREC exhibits superior speed and energy efficiency (as will be discussed in the following section and shown in Table 3), as well as a significant improvement in memory efficiency. This is largely due to the fact that RISE computes the auto-correlation function, with the default setting being 100 auto-correlation terms, which inherently makes SIREC more suited for local operation on micro-controllers for classification tasks. Furthermore, SIREC tends to demonstrate marginally better average accuracy. To support this claim, Figure 6, as an extension to Figure 5, provides a detailed depiction of the F1 scores between RISE and SIREC over the 100 iterations.
@@
A significant advantage of the Random Interval Generation concept for feature extraction, as implemented in SIREC, is its inherent reduction or elimination of the need for expert knowledge about the collected data to train the model. For instance, a stochastic search across various randomly generated intervals within specified minimum and maximum lengths allows for the bypassing of such expert knowledge. This stochastic search method ensures the identification of the ideal data segment for feature extraction without necessitating extensive manual data analysis. Moreover, since SIREC is based on DTs, which represent a symbolic form of representation, acquiring expert knowledge about the data post-training is straightforward. Our stochastic search for optimal SIREC parameters (thus, the best random intervals) has achieved accuracies up to 100% in some cases. The results of this stochastic search, along with the parameters used for SIREC, are presented in Table 2.
@@ 3.2 ENERGY MEASUREMENTS
In the analysis of energy consumption, our investigation reveals significant insights into the trade-offs between computational energy efficiency and predictive accuracy across three distinct algorithms: our proposed algorithm (SIREC), Time Series Forest Classifier (TSFC)[35] and RISE. The measurement method and experiment setup is based upon our previous work [36, 37] and explained in detail by [38]. The resource measurements adhere to the methodology and guidelines outlined in the Green Software Measurement Model [39]. Measurements were carried out using a computer system that included an Intel Core i5-650 CPU. The system’s memory was composed of two 2 GB RAM modules, totaling 4 GB. For storage, the computer utilized a dual-setup, comprising a 500 GB hard disk drive (HDD) and a 250 GB solid-state drive (SSD). The analysis is structured around various scenarios–configurations of segment lengths and numbers of estimators, focusing on key metrics such as energy consumption during training and testing phases, computational time, CPU usage, and F1 score. These scenarios, along with their corresponding results listed in Table 3, were tested on the same datasets, $T_{1}$ and $T_{2}$ for training, and $T_{3}$ and $T_{4}$ for testing, as detailed in Table 1. It is important to highlight that a random state parameter was employed across all algorithms to ensure comparability and control over experimental variability. A value of 7 was selected arbitrarily for the random state, with no intent to favor any specific algorithm, especially not SIREC.
@@ 4 CONCLUSION
5 ACKNOWLEDGMENT
The authors would like to thank the students, Mr. Marvin Schacht for his support in constructing the demonstrators, and Mr. Marcel Andres for his assistance with the energy measurements.